# MEGADance: Mixture-of-Experts Architecture for Genre-Aware 3D Dance Generation

### v2

@@ 1 Introduction
Music-to-dance generation is a crucial task that translates auditory input into dynamic motion, with significant applications in virtual reality, choreography, and digital entertainmentli2021ai ; li2023finedance . By automating this process, it enables deeper exploration of the intrinsic relationship between music and movementyang2024beatdance , while expanding possibilities for creative content generation. Due to its broad impact, music-to-dance generation has attracted increasing attentionsiyao2022bailando ; siyao2023bailando++ ; li2024lodge++ .
Current music-to-dance generation approaches have witnessed rapid progress and can be broadly categorized into two paradigmssiyao2022bailando ; zhuang2023gtn : (1) One-stage methods directly map musical features to human motionli2024lodge ; li2024lodge++ ; huang2020dance . (2) Two-stage methods first construct choreographic units and then learn their probability distributions conditioned on music siyao2022bailando ; siyao2023bailando++ ; siyao2024duolando . However, previous methods only treat genre as a weak auxiliary bias rather than the core semantic driverli2024exploring ; li2023finedance ; li2024lodge , facing several essential problems such as misaligned music-motion synchronization and disrupted dance genre continuity. For example, A Uyghur movements clip is inappropriately mixed into the Popping routine, when a typical Popping music exhibits complex transitions in rhythm and intensity, as shown in Fig. 1. This oversight leads to unsatisfactory visual effects, and fails to meet genre-specific user demands.
@@
To address these limitations, we propose MEGADance, the first Mixture-of-Expertsshazeer2017outrageously (MoE) architecture for Genre-Aware 3D Dance Generation. By decoupling choreographic consistency into dance generality, modeled by Universal Experts shared across all genres, and genre specificity, captured by Specialized Experts selected via genre-guided hard routing, MEGADance enables robust musical alignment and fine-grained stylistic fidelity. Due to its structured inductive bias, MEGADance exhibits strong genre control and remains robust even in the presence of modality conflicts, such as generating a Breaking dance for soft-paced Chinese music while preserving rhythmic and dynamic alignment.
@@
Specifically, MEGADance comprises two stages. (1) High-Fidelity Dance Quantization Stage (HFDQ), which encodes dance motions into a latent representation. In the HFDQ stage, we introduce Finite Scalar Quantizationmentzer2023finite (FSQ), which replaces traditional VQ-VAE codebooks to mitigate codebook collapse and enhance latent diversity. Additionally, we impose kinematic constraints by simultaneously reconstructing 3D joints through Forward Kinematicsloper2023smpl from SMPLloper2023smpl , and dynamic constraints by reconstructing dances while considering velocity and acceleration, to enhance spatio-temporal coherence. (2) Genre-Aware Dance Generation Stage (GADG), which maps music into the latent representation. In the GADG, the Universal Experts and Specialized Experts work synergistically to jointly modeling dance generality and genre specificity. For genre-disentangled expert design, Universal Experts model universal rhythmic and temporal structures across all genres, enhancing robustness to diverse musical inputs and stabilizing cross-modal alignment; and Specialized Experts specialize in fine-grained stylistic variations unique to each genre, guided by hard routing to effectively disentangle genre-dependent features from genre-invariant dynamics. For expert structure, each expert, whether Specialized or Universal, adopts an autoregressive Mamba-Transformer hybrid backbone, combining Mamba’s efficient intra-modal local dependency capture gu2023mamba with the Transformer’s cross-modal global contextual understandingvaswani2017attention , thereby enabling the generation of temporally coherent and musically aligned dance motions.
@@
The contributions of our work can be summarized as: (1) We introduce MEGADance, the first Mixture-of-Experts (MoE) architecture for music-to-dance generation, designed to enhance choreographic consistency by decoupling it into dance generality and genre specificity. MEGADance achieves state-of-the-art (SOTA) performance and demonstrates robust genre controllability, as demonstrated through extensive qualitative and quantitative experiments on the AIST++li2021ai and FineDanceli2023finedance datasets. (2) We propose a High-Fidelity Dance Quantization framework that introduces FSQ with kinematic-dynamic dual constraints, ensuring complete codebook utilization (100% vs. VQ-VAEoord2017neural ’s 75%) while achieving excellent reconstruction accuracy. (3) We design a Mamba-Transformer hybrid backbone for music-to-dance generation, combining Mamba’s efficient intra-modal local dependency capture with the Transformer’s cross-modal global contextual understanding.
@@ 2.1 One-Stage Music-to-Dance Generation
Music and dance are deeply interconnected, leading to significant advancements in the field of music-driven 3D dance generation. Researchers utilize musical features extracted via tools like Librosali2021ai , Jukeboxtseng2023edge , and MERTyang2024codancers to predict human motion, including SMPLloper2023smpl parametersli2021ai and body keypointssiyao2022bailando . Early methods primarily employ encoder-decoder architectures to directly obtain entire human motion sequence lee2018listen ; tang2018dance ; le2023music ; correia2024music ; li2021ai . Recognizing the natural hierarchical structure of human joints, some researchers introduced Graph Convolutional Networks (GCNs)ferreira2021learning ; yan2019convolutional to enhance interaction at the joint level, thereby improving the biomedical plausibility of the generated motions. In AIGC, Generative Adversarial Networks (GANs) are widely applied, some researchers introduced it in music-to-dance taskschen2021choreomaster ; huang2020dance ; yang2024cohedancers . Specifically, GANs’ generators produce dance motions from music, with discriminators providing feedback to guide generated motions more natural. Recently, Diffusion Models have shown remarkable success in various AIGC tasks, with notable applications extending to the music-to-dance domaintseng2023edge ; li2023finedance ; li2024lodge ; li2024lodge++ ; le2023controllable , but the computational cost of the sampling process remains high, especially in long-sequence generation scenarios for the music-to-dance task. However, the lack of explicit constraints to maintain the generated pose within proper spatial boundaries often leads to nonstandard poses that extend beyond the dancing subspace during inference, resulting in low dance quality in practical applications.
@@ 2.2 Two-Stage Music-to-Dance Generation
Dance Quantization Stage. Traditional methods ye2020choreonet ; aristidou2022rhythm ; huang2022genre ; chen2021choreomaster construct choreographic units through uniform segmentation of motion sequence, incurring high computational overhead. Recent works gong2023tm2d ; guo2022generating employ VQ-VAE for intelligent unit construction, significantly reducing time/space complexity. Considering the relative independence of upper and lower body movements, siyao2022bailando ; siyao2023bailando++ construct choreographic unit for lower and upper parts separately, improving motion reconstruction through expanded unit capacity ($L\to L\times L$). However, above works predominantly operate on 3D human body keypoints, which lack expressiveness in capturing nuanced motion details. siyao2023bailando++ ; li2024lodge++ construct choreographic units in the SMPL pose space but apply uniform treatment across joints, neglecting the body’s kinematic hierarchy, such as root errors propagate globally through kinematic chains while hand errors remain localized.
Dance Generation Stage. To model choreographic unit distributions, Choreomaster chen2021choreomaster employs a GRU-based backbone, while DanceRevolution huang2020dance uses RNNs. Recent works like Bailando siyao2022bailando and Bailando++siyao2023bailando++ adopt cross-modal Transformers for improved temporal modeling and music-motion alignment. Moreover, Everything2Motion fan2024everything2motion and TM2D gong2023tm2d leverage pretrained models in text-to-motionguo2022generating ; guo2022tm2t generation to improve motion quality, often at the expense of choreographic complexity and creativity. To enrich input representations, li2023finedance ; li2024lodge ; huang2022genre ; zhuang2023gtn introduce genre information via shallow fusion, such as cross attention huang2022genre or feature addition zhuang2023gtn . However, these approaches remain insufficient for achieving robust genre controllability, particularly under cross-modal conflicts, such as generating Breaking dances conditioned on slow-tempo traditional Chinese music.
@@ 3.1 Problem Definition
Given a music sequence $M=\{m_{0},m_{1},...,m_{T}\}$ and a dance genre label $g$, our objective is to synthesize the corresponding dance sequence $S=\{s_{0},s_{1},...,s_{T}\}$, where $m_{t}$ and $s_{t}$ denotes the music and dance feature at time step $t$. We define each music feature $m_{t}$ as a 35-dim vectorli2024lodge extracted by Librosamcfee2015librosa , including 20-dim MFCC, 12-dim Chroma, 1-dim Peak, 1-dim Beat and 1-dim Envelope. We encode genre label $g$ as a one-hot vector. We represent each dance feature as a 147-dim vector $s_{t}=[\tau;\theta]$, where $\tau$ and $\theta$ encapsulate the root translation and 6-dim rotation representationzhou2019continuity of the SMPLloper2023smpl model, respectively. Furthermore, we synchronize the music sequence with the dance sequence at a temporal granularity of 30 frames per second.
@@ 3.2 High-Fidelity Dance Quantization
Our 3D motion reconstruction approach, illustrated in Fig. 2, initiates with a Dance Encoder $\mathbf{E}$ (a three-layer 1D-CNN for information aggregation and a two-layer MLP for dimension adjustment) encoding the dance sequence $S=\{S^{u},S^{l}\}$ into context-aware features $\mathbf{z}=\{\mathbf{z^{u}},\mathbf{z^{l}}\}$. These features are quantized using Finite Scalar Quantization (FSQ) to obtain $\hat{\mathbf{z}}=\{\hat{\mathbf{z^{u}}},\hat{\mathbf{z^{l}}}\}$, which are then decoded by Dance Decoder $\mathbf{D}$ (a two-layer MLP for dimension adjustment and a three-layer 1D TransConv for information restoration) to reconstruct the dance movement $\hat{S}=\{\hat{S^{u}},\hat{S^{l}}\}$. To resolve the codebook collapse problem caused by the conventional VQ-VAEvan2017neural based quantization, we adopt FSQ. By replacing the discrete "argmin" codebook selection with scalar quantization via differentiable bounded rounding, FSQ enables balanced utilization and stable gradient propagation:
@@
where $f(\cdot)$ is the bounding function, setting as the $\mathrm{sigmoid}(\cdot)$ function in our practice. Each channel in $\hat{\mathbf{z}}$ will be quantized into one of the unique $L$ integers, therefore we have $\hat{\mathbf{z}}\in\{1,\dots,L\}^{d}$. The codebook size is calculated as $k=\prod_{i=1}^{d}L_{i}$, and $L,d$ are super parameter. In conclusion, FSQ with Motion Decomposition expands its effective motion representation capacity, thereby enhancing the diversity of subsequent generated dance.
@@
The $\mathcal{L}_{\text{smpl}}$ is the reconstruction loss that ensures the predicted 3D SMPL sequence closely aligns with the ground truth. Simple reconstruction on SMPL parameters treats all joints equally, neglecting the complex hierarchical tree structure of human body joints, different joints vary in their tolerance to errors. For instance, errors at the root node propagate throughout all nodes, whereas errors at the hand node primarily affect only itself. Thus, we execute Forward Kineticloper2023smpl techniques to derive 3D joints and apply reconstruction constraints $\mathcal{L}_{\text{joint}}$ between $\hat{J}$ and $J$. Moreover, the reconstruction loss accounts not only for the spatial positions but also for the velocities ($\alpha_{1}$) and accelerations ($\alpha_{2}$) of the movements:
@@ 3.3.1 Mixture-of-Experts Architecture
As illustrated in Fig. 2, we perform cross-modal autoregressive generation. Given music features $m_{1:T}$ extracted using Librosa, dance genre label $g$, and the previous pose codes $p_{0:T-1}=\{p_{0:T-1}^{l},p_{0:T-1}^{u}\}$ encoded by the Motion Tokenization $\mathbf{E}$ in HFDQ, the GADG predicts action probabilities $a_{0:T-1}=\{a_{0:T-1}^{l},a_{0:T-1}^{u}\}$ of every $z_{i}\in\mathcal{Z}$, using an $L$-layer MoE architecture. To align the predicted action probabilities $a_{0:T-1}$ with the next pose codes $p_{1:T}$, we employ a supervised Cross-Entropysiyao2022bailando loss, where each predicted action $a_{t}$ is matched to its corresponding target pose code $p_{t+1}$. The inference of GADG includes: 1) Short sequences ($\leq 5.5s$) via autoregressive generation, 2) Long sequences via sliding-window prediction with 5.5s overlap.
@@ 3.3.2 Genre-Disentangled Experts Design
Specialized Experts. The Specialized Experts are designed to capture genre-specific stylistic patterns, motivated by two core considerations: (1) Structural Inductive Bias: By isolating parameters across experts, the model enforces separation of genre-specific motion motifs (e.g., Krump’s grounded explosiveness vs. Contemporary’s fluid transitions), thereby preserving distinct stylistic representations. This separation also introduces genre-aware control priors that mitigate cross-genre interference, which is critical for genre-conditioned dance generation. (2) Computational Efficiency: Leveraging sparse MoE design Shazeer2017 , each input is routed to a single expert, significantly reducing parameter redundancy and computational cost.
Universal Experts. The Universal Expert learns generalizable representations to complement Specialized Experts through two key roles: (1) Fundamental Choreographic Prior: It learns shared low-level patterns across genres (e.g. periodicity, beat synchronization, and biomechanical consistency). In contrast, models relying solely on Specialized Experts often fail under modality mismatch (e.g., producing static or repetitive movements when Ballet music is processed by a Popping Expert). The Universal Expert provides a genre-agnostic prior that enhances stability and expressiveness under complex input conditions. (2) Disentangled Representation: By disentangling shared and genre-specific factors, the model allows each expert to specialize in distinct subspaces, enhancing generation qualitylocatello2020weakly .

### v3

@@ 1 Introduction
Music-to-dance generation is a crucial task that translates auditory input into dynamic motion, with significant applications in virtual reality, choreography, and digital entertainmentLi et al. (2021, 2023). By automating this process, it enables deeper exploration of the intrinsic relationship between music and movementYang et al. (2024c), while expanding possibilities for creative content generation. Due to its broad impact, music-to-dance generation has attracted increasing attentionSiyao et al. (2022, 2023); Li et al. (2024b).
Current music-to-dance generation approaches have witnessed rapid progress and can be broadly categorized into two paradigmsSiyao et al. (2022); Zhuang et al. (2023): (1) One-stage methods directly map musical features to human motionLi et al. (2024c, b); Huang et al. (2020). (2) Two-stage methods first construct choreographic units and then learn their probability distributions conditioned on music Siyao et al. (2022, 2023, 2024). However, previous methods only treat genre as a weak auxiliary bias rather than the core semantic driverLi et al. (2024a, 2023, c), facing several essential problems such as misaligned music-motion synchronization and disrupted dance genre continuity. For example, A Uyghur movements clip is inappropriately mixed into the Popping routine, when a typical Popping music exhibits complex transitions in rhythm and intensity, as shown in Fig. 1. This oversight leads to unsatisfactory visual effects, and fails to meet genre-specific user demands.
@@
To address these limitations, we propose MEGADance, the first Mixture-of-ExpertsShazeer et al. (2017a) (MoE) architecture for Genre-Aware 3D Dance Generation. By decoupling choreographic consistency into dance generality, modeled by Universal Experts shared across all genres, and genre specificity, captured by Specialized Experts selected via genre-guided hard routing, MEGADance enables robust musical alignment and fine-grained stylistic fidelity. Due to its structured inductive bias, MEGADance exhibits strong genre control and remains robust even in the presence of modality conflicts, such as generating a Breaking dance for soft-paced Chinese music while preserving rhythmic and dynamic alignment.
@@
Specifically, MEGADance comprises two stages. (1) High-Fidelity Dance Quantization Stage (HFDQ), which encodes dance motions into a latent representation. In the HFDQ stage, we introduce Finite Scalar QuantizationMentzer et al. (2023) (FSQ), which replaces traditional VQ-VAE codebooks to mitigate codebook collapse and enhance latent diversity. Additionally, we impose kinematic constraints by simultaneously reconstructing 3D joints through Forward KinematicsLoper et al. (2023) from SMPLLoper et al. (2023), and dynamic constraints by reconstructing dances while considering velocity and acceleration, to enhance spatio-temporal coherence. (2) Genre-Aware Dance Generation Stage (GADG), which maps music into the latent representation. In the GADG, the Universal Experts and Specialized Experts work synergistically to jointly modeling dance generality and genre specificity. For genre-disentangled expert design, Universal Experts model universal rhythmic and temporal structures across all genres, enhancing robustness to diverse musical inputs and stabilizing cross-modal alignment; and Specialized Experts specialize in fine-grained stylistic variations unique to each genre, guided by hard routing to effectively disentangle genre-dependent features from genre-invariant dynamics. For expert structure, each expert, whether Specialized or Universal, adopts an autoregressive Mamba-Transformer hybrid backbone, combining Mamba’s efficient intra-modal local dependency capture Gu and Dao (2023) with the Transformer’s cross-modal global contextual understandingVaswani (2017), thereby enabling the generation of temporally coherent and musically aligned dance motions.
@@
The contributions of our work can be summarized as: (1) We introduce MEGADance, the first Mixture-of-Experts (MoE) architecture for music-to-dance generation, designed to enhance choreographic consistency by decoupling it into dance generality and genre specificity. MEGADance achieves state-of-the-art (SOTA) performance and demonstrates robust genre controllability, as demonstrated through extensive qualitative and quantitative experiments on the AIST++Li et al. (2021) and FineDanceLi et al. (2023) datasets. (2) We propose a High-Fidelity Dance Quantization framework that introduces FSQ with kinematic-dynamic dual constraints, ensuring complete codebook utilization (100% vs. VQ-VAEvan den Oord et al. (2017)’s 75%) while achieving excellent reconstruction accuracy. (3) We design a Mamba-Transformer hybrid backbone for music-to-dance generation, combining Mamba’s efficient intra-modal local dependency capture with the Transformer’s cross-modal global contextual understanding.
@@ 2.1 One-Stage Music-to-Dance Generation
Music and dance are deeply interconnected, leading to significant advancements in the field of music-driven 3D dance generation. Researchers utilize musical features extracted via tools like LibrosaLi et al. (2021), JukeboxTseng et al. (2023), and MERTYang et al. (2024a) to predict human motion, including SMPLLoper et al. (2023) parametersLi et al. (2021) and body keypointsSiyao et al. (2022). Early methods primarily employ encoder-decoder architectures to directly obtain entire human motion sequence Lee et al. (2018); Tang et al. (2018); Le et al. (2023b); Correia and Alexandre (2024); Li et al. (2021). Recognizing the natural hierarchical structure of human joints, some researchers introduced Graph Convolutional Networks (GCNs)Ferreira et al. (2021); Yan et al. (2019) to enhance interaction at the joint level, thereby improving the biomedical plausibility of the generated motions. In AIGC, Generative Adversarial Networks (GANs) are widely applied, some researchers introduced it in music-to-dance tasksChen et al. (2021); Huang et al. (2020); Yang et al. (2024b). Specifically, GANs’ generators produce dance motions from music, with discriminators providing feedback to guide generated motions more natural. Recently, Diffusion Models have shown remarkable success in various AIGC tasks, with notable applications extending to the music-to-dance domainTseng et al. (2023); Li et al. (2023, 2024c, 2024b); Le et al. (2023a), but the computational cost of the sampling process remains high, especially in long-sequence generation scenarios for the music-to-dance task. However, the lack of explicit constraints to maintain the generated pose within proper spatial boundaries often leads to nonstandard poses that extend beyond the dancing subspace during inference, resulting in low dance quality in practical applications.
@@ 2.2 Two-Stage Music-to-Dance Generation
Dance Quantization Stage. Traditional methods Ye et al. (2020); Aristidou et al. (2022); Huang et al. (2022); Chen et al. (2021) construct choreographic units through uniform segmentation of motion sequence, incurring high computational overhead. Recent works Gong et al. (2023); Guo et al. (2022a) employ VQ-VAE for intelligent unit construction, significantly reducing time/space complexity. Considering the relative independence of upper and lower body movements, Siyao et al. (2022, 2023) construct choreographic unit for lower and upper parts separately, improving motion reconstruction through expanded unit capacity ($L\to L\times L$). However, above works predominantly operate on 3D human body keypoints, which lack expressiveness in capturing nuanced motion details. Siyao et al. (2023); Li et al. (2024b) construct choreographic units in the SMPL pose space but apply uniform treatment across joints, neglecting the body’s kinematic hierarchy, such as root errors propagate globally through kinematic chains while hand errors remain localized.
Dance Generation Stage. To model choreographic unit distributions, Choreomaster Chen et al. (2021) employs a GRU-based backbone, while DanceRevolution Huang et al. (2020) uses RNNs. Recent works like Bailando Siyao et al. (2022) and Bailando++Siyao et al. (2023) adopt cross-modal Transformers for improved temporal modeling and music-motion alignment. Moreover, Everything2Motion Fan et al. (2024) and TM2D Gong et al. (2023) leverage pretrained models in text-to-motionGuo et al. (2022a, b) generation to improve motion quality, often at the expense of choreographic complexity and creativity. To enrich input representations, Li et al. (2023, 2024c); Huang et al. (2022); Zhuang et al. (2023) introduce genre information via shallow fusion, such as cross attention Huang et al. (2022) or feature addition Zhuang et al. (2023). However, these approaches remain insufficient for achieving robust genre controllability, particularly under cross-modal conflicts, such as generating Breaking dances conditioned on slow-tempo traditional Chinese music.
@@ 3.1 Problem Definition
Given a music sequence $M=\{m_{0},m_{1},...,m_{T}\}$ and a dance genre label $g$, our objective is to synthesize the corresponding dance sequence $S=\{s_{0},s_{1},...,s_{T}\}$, where $m_{t}$ and $s_{t}$ denotes the music and dance feature at time step $t$. We define each music feature $m_{t}$ as a 35-dim vectorLi et al. (2024c) extracted by LibrosaMcFee et al. (2015), including 20-dim MFCC, 12-dim Chroma, 1-dim Peak, 1-dim Beat and 1-dim Envelope. We encode genre label $g$ as a one-hot vector. We represent each dance feature as a 147-dim vector $s_{t}=[\tau;\theta]$, where $\tau$ and $\theta$ encapsulate the root translation and 6-dim rotation representationZhou et al. (2019) of the SMPLLoper et al. (2023) model, respectively. Furthermore, we synchronize the music sequence with the dance sequence at a temporal granularity of 30 frames per second.
@@ 3.2 High-Fidelity Dance Quantization
Our 3D motion reconstruction approach, illustrated in Fig. 2, initiates with a Dance Encoder $\mathbf{E}$ (a three-layer 1D-CNN for information aggregation and a two-layer MLP for dimension adjustment) encoding the dance sequence $S=\{S^{u},S^{l}\}$ into context-aware features $\mathbf{z}=\{\mathbf{z^{u}},\mathbf{z^{l}}\}$. These features are quantized using Finite Scalar Quantization (FSQ) to obtain $\hat{\mathbf{z}}=\{\hat{\mathbf{z^{u}}},\hat{\mathbf{z^{l}}}\}$, which are then decoded by Dance Decoder $\mathbf{D}$ (a two-layer MLP for dimension adjustment and a three-layer 1D TransConv for information restoration) to reconstruct the dance movement $\hat{S}=\{\hat{S^{u}},\hat{S^{l}}\}$. To resolve the codebook collapse problem caused by the conventional VQ-VAEVan Den Oord et al. (2017) based quantization, we adopt FSQ. By replacing the discrete "argmin" codebook selection with scalar quantization via differentiable bounded rounding, FSQ enables balanced utilization and stable gradient propagation:
@@
[…]
@@
The $\mathcal{L}_{\text{smpl}}$ is the reconstruction loss that ensures the predicted 3D SMPL sequence closely aligns with the ground truth. Simple reconstruction on SMPL parameters treats all joints equally, neglecting the complex hierarchical tree structure of human body joints, different joints vary in their tolerance to errors. For instance, errors at the root node propagate throughout all nodes, whereas errors at the hand node primarily affect only itself. Thus, we execute Forward KineticLoper et al. (2023) techniques to derive 3D joints and apply reconstruction constraints $\mathcal{L}_{\text{joint}}$ between $\hat{J}$ and $J$. Moreover, the reconstruction loss accounts not only for the spatial positions but also for the velocities ($\alpha_{1}$) and accelerations ($\alpha_{2}$) of the movements:
@@ 3.3.1 Mixture-of-Experts Architecture
As illustrated in Fig. 2, we perform cross-modal autoregressive generation. Given music features $m_{1:T}$ extracted using Librosa, dance genre label $g$, and the previous pose codes $p_{0:T-1}=\{p_{0:T-1}^{l},p_{0:T-1}^{u}\}$ encoded by the Motion Tokenization $\mathbf{E}$ in HFDQ, the GADG predicts action probabilities $a_{0:T-1}=\{a_{0:T-1}^{l},a_{0:T-1}^{u}\}$ of every $z_{i}\in\mathcal{Z}$, using an $L$-layer MoE architecture. To align the predicted action probabilities $a_{0:T-1}$ with the next pose codes $p_{1:T}$, we employ a supervised Cross-EntropySiyao et al. (2022) loss, where each predicted action $a_{t}$ is matched to its corresponding target pose code $p_{t+1}$. The inference of GADG includes: 1) Short sequences ($\leq 5.5s$) via autoregressive generation, 2) Long sequences via sliding-window prediction with 5.5s overlap.
@@ 3.3.2 Genre-Disentangled Experts Design
Specialized Experts. The Specialized Experts are designed to capture genre-specific stylistic patterns, motivated by two core considerations: (1) Structural Inductive Bias: By isolating parameters across experts, the model enforces separation of genre-specific motion motifs (e.g., Krump’s grounded explosiveness vs. Contemporary’s fluid transitions), thereby preserving distinct stylistic representations. This separation also introduces genre-aware control priors that mitigate cross-genre interference, which is critical for genre-conditioned dance generation. (2) Computational Efficiency: Leveraging sparse MoE design Shazeer et al. (2017b), each input is routed to a single expert, significantly reducing parameter redundancy and computational cost.
Universal Experts. The Universal Expert learns generalizable representations to complement Specialized Experts through two key roles: (1) Fundamental Choreographic Prior: It learns shared low-level patterns across genres (e.g. periodicity, beat synchronization, and biomechanical consistency). In contrast, models relying solely on Specialized Experts often fail under modality mismatch (e.g., producing static or repetitive movements when Ballet music is processed by a Popping Expert). The Universal Expert provides a genre-agnostic prior that enhances stability and expressiveness under complex input conditions. (2) Disentangled Representation: By disentangling shared and genre-specific factors, the model allows each expert to specialize in distinct subspaces, enhancing generation qualityLocatello et al. (2020).